OpenStudy (anonymous):

Simplify. Express the quotient as a radical expression. ^11SQRTX^10 / ^11SQRTx^6 the options are A) ^0 SQRT x^4 B) ^11 SQRT x^16 C) ^11 SQRT x^4 D) ^121 SQRT x^60

OpenStudy (dumbcow):

\[\frac{\sqrt[11]{x^{10}}}{\sqrt[11]{x^{6}}} = \frac{x^{10/11}}{x^{6/11}} = x^{10/11 - 6/11} = x^{4/11} = \sqrt[11]{x^{4}}\]